# and quantum information science

### v2

@@ 5.2 Geometric features of Nielsen complexity
Returning to the cost function (25), consider taking $p=1$. When $p=1$, $F_{p{=}1,\vec{q}}[Y^{I}]$ has a natural physical interpretation: it counts (with an appropriate measure) the number of gates used to build the continuous trajectory in the group manifold SU($N$). However, $F_{p{=}1,\vec{q}}[Y^{I}]$ has the disadvantage of not being smooth. We therefore cannot apply the calculus of variations to the geodesics. When $p=2$, the norm (25) induces a Riemannian metric on the group manifold, simplifying the study of the geodesics. In this case, the map (22) is a Riemannian submersion [101] (defined in [102]). Heuristically, a Riemannian submersion is a smooth mapping from a higher-dimensional manifold to a lower-dimensional one. This mapping preserves distances in the directions perpendicular to the fibers. [^10] The fibers are points in the higher-dimensional space that are mapped to the same point in the lower-dimensional manifold. In our setting, the fibers correspond to the maximal subgroup $\mathrm{SU}(N-1)\times\mathrm{U}(1)$. The submersion maps $\mathrm{SU}(N)$ to $\mathbb{CP}^{N-1}$. These facts provide a systematic means of determining the metric induced by the cost-function minimization (23) on the state space. The Nielsen complexity literature has mainly focused on $p=1,2$ [99, 100, 103, 104, 105, 106, 107, 108, 109, 110, 55, 40, 111, 112, 101, 42, 113, 97, 114, 19, 44, 115].
One may wonder whether Nielsen’s complexity is related to discrete gate complexity. The answer is yes: Nielsen’s complexity provides both upper and lower bounds on the gate complexity, provided the cost function satisfies certain conditions [93, 94]. The authors of [93, 94] further show that, by sufficiently penalizing the generators acting on several qubits, one can force the geodesic trajectory that reaches a unitary in the complexity geometry with arbitrary precision to only use generators acting on one or two qubits. The penalties required for this construction typically make the manifold highly curved. However, as we will discuss, Nielsen’s complexity is universal at long distances within a certain equivalence class of metrics, and this equivalence class can be shown to also contain moderately-curved metrics. A priori, this fact makes the computation of Nielsen’s complexity easier by using tools from differential geometry. Finally, let us mention that when the motion in the unitary manifold is restricted to a small subgroup, exact or semi-analytic results can be obtained, which we discuss in later subsections.
@@
There is substantial freedom in choosing a cost function. However, there is a large universality class of cost functions leading to metrics on the group manifold that may vary significantly at short distances, but differ at large distances only polynomially in the path length and in the physical system’s size [97, 19]. To simplify the rest of this subsection, we will take the penalty factors to only depend on the number of qubits $k$ on which a gate acts, and denote them as $q_{k}$. Sometimes $k$ is referred to as the weight of the gate – $q_{k}$ should not be confused with the $q_{I}$ introduced in Eq. (25). The notation $q_{k}$ means that generators $T_{I}$ that have the same weight $k$ share the same penalty factor.
Consider increasing any penalty factor while keeping the others constant. Above a critical value $\bar{q}_{k}$, the complexity ceases to depend on that penalty factor because we can implement this generator with weight $k$ as a composition of other, cheaper gates. The reason is as follows. Consider any fixed $m\geq 2$. Using the $m$-local gates, one can reconstruct all the $m^{\prime}$-local gates for every $m^{\prime}$ [15]. Consequently, one can identify a certain critical schedule $\{\bar{q}_{k}\}$ in terms of which we can define a universality class of cost functions that yield metrics with the same long distance properties [97]. Specifically, given the critical schedule, we consider penalty factors $\{q_{k}\}$ satisfying $q_{2}=\bar{q}_{2}$ and, for all $m>2$, $q_{m}\geq\bar{q}_{m}$. Metrics arising from this universality class of cost functions may differ greatly at short separations but lead to approximately equal distances at long separations. That is, short and long distance scales decouple. The critical schedule is conjectured to be the only universality class member for which the conjugate points are pushed to infinity, and geodesics are straight lines for an exponential time in the number of qubits [97]. [^11] A conjugate point occurs along a geodesic if there is a local shortcut from the start of the trajectory to some point along it (the notion of local shortcut will be more precisely defined below). Such conjugate points can be diagnosed by locally perturbing a geodesic. If the perturbation does not change the length, the geodesic can be deformed to find a shorter geodesic between the same endpoints. There can also be global obstructions to minimality of a geodesic, such as geodesics encircling a sphere in opposite directions along a diameter. We will discuss conjugate points in more detail in Sec. 5.4. Brown argued that the same universality class consists of metrics whose associated Nielsen complexity, measured by the geodesic length from the origin, is polynomially equivalent to the gate complexity. At the same time, it is possible to pick a metric inside the universality class such that the unitary manifold has modest curvature [19], thus making computations easier.
@@ 5.3 Constraints on the cost function from expected time dynamics
In Sec. 4.1, we showed that quantum complexity for discrete time circuits exhibits two behaviors: linear growth which can last for an exponentially long time, and the switchback effect. These derivations relied on simple counting arguments. Afterwards, we quantitatively demonstrated this linear growth for random unitary circuits, by using algebraic geometry and unitary designs (Sections 4.34.6). We would like to understand which complexity geometries reproduce the above key features of complexity when applied to the unitary time evolution generated by a chaotic Hamiltonian. To do so, we must select an appropriate metric on the $\mathrm{SU}(N)$ manifold. The requirements turn out to be [95, 50, 55]:
@@
    The group manifold must have a negative average sectional curvature [95, 50], such that nearby geodesics diverge. This is necessary, but not sufficient, for establishing the above key features for chaotic and ergodic Hamiltonians.
@@
Let us comment on the implications of these requirements. First, consider Eq. (23) which yields a metric on the space of states rather than on unitaries. Do the above properties have a reflection in the induced geometry on the space of states? In fact the induced metric is not homogeneous, complicating the study of the geodesics and curvature in the space of states [108]. Nonetheless, as mentioned earlier, for the case of Riemannian cost functions, the map (22) is a Riemannian submersion [101] (defined in [102]). Riemannian submersions relate sectional curvatures in the space of unitaries and sectional curvatures in the space of states, via the O’Neill formula [116]. The latter implies that the sectional curvature along a plane in the space of states is lower-bounded by the sectional curvature along an appropriate plane in the unitary manifold.
@@
Second, the group manifold equipped with the Cartan-Killing metric—the standard metric on SU$(N)$—does not satisfy the above requirements because it has only constant positive sectional curvatures. How can we construct a geometry that has negative sectional curvatures? We can assign the penalty factors such that some of the $\mathfrak{su}(N)$ algebra’s nonvanishing commutators have the structure $[easy,\,easy]=\,hard$. The easy and hard refer to generators associated with small and large penalty factors in the cost function. When the penalty factors are associated to generators such that the commutators take the above-mentioned form, then the Pythagorean theorem for curved space guarantees some negative sectional curvatures [108]. The latter statement was proven in the contexts of the unitary manifolds of qubits and qutrits [108, 101]. [^12] A qudit is a $d$-level quantum system, represented by a Hilbert space $\mathbb{C}^{d}$. A qubit has $d=2$; and a qutrit, $d=3$. Note that manifolds that are equipped with right-invariant metrics, like the ones we are interested in, must also have some positive sectional curvatures; otherwise, they are flat [117].
@@
References [112, 114] studied time-dependent trajectories generated by the SYK Hamiltonian on the unitary manifold. The authors showed that the geodesic distance from the origin to the time evolution operator is lower-bounded by a geodesic whose length grows linearly with time. This linear growth is truncated at conjugate points along the geodesic. There is a shorter path connecting the origin to a conjugate point than the one produced by extending the earlier time geodesic. It is possible that the linear growth is truncated even earlier by global obstructions – geodesics that wind in some other direction around the unitary group to the same point – but these are much more difficult to study, as we will discuss in Sec. 5.4. Conjugate points in the critical complexity geometry described at the end of Sec. 5.2 are pushed to infinity [97]. Therefore, the linear growth of Nielsen’s complexity in that geometry persists for a time exponential in the number of qubits. In the context of spin chains, Brown showed that many complexity geometries have diameters that are exponentially large in the number of qubits [19]. This exponential relation is necessary for complexity to saturate at an exponentially large value (Fig. 3).
@@
We have argued that Nielsen complexity exhibits one behavior characteristic of quantum complexity: the initial linear growth. Nielsen complexity also reproduces the other behavior—the switchback effect—in toy models [55]. Qualitative features similar to the switchback effect are also displayed by single-qubit and two-qubit systems [118]. However, a comprehensive picture of the emergence of the switchback effect from the complexity geometry is still lacking and is an open question for future research.
@@ 5.4 Physical criteria for complexity growth: general results and applications
How is Nielsen complexity related to standard concepts in physics, like energy spectra, correlation functions, and Schrödinger dynamics? Several connections were uncovered by the authors of [112, 114], who investigated physical criteria that control the length of time over which Nielsen complexity will grow during Hamiltonian time evolution. We will discuss their results below, and use this discussion as a way of illustrating how to perform practical calculations of Nielsen complexity.
@@
For concreteness, let us consider a time-independent Hamiltonian acting on an $N$–dimensional Hilbert space. Time evolution is then governed by the unitary operator $U(t)=e^{-iHt}$. We want to think of this operator as tracing a trajectory on the $\mathrm{SU}(N)$ group manifold from the identity $\mathbbm{1}$ at $t=0$ to $U(t)$. Following Nielsen, we will regard the complexity of $U(t)$ as measured by the length of the shortest geodesic from $\mathbbm{1}$ to $U(t)$ in a certain complexity metric that we will define below.
@@
In fact, finding the shortest geodesic on $\mathrm{SU}(N)$ for general $N$ is a difficult problem in any metric because the group manifold has a complicated fiber bundle topology. So, if we find one geodesic between two group elements by solving the appropriate differential equation, there may well be others that start at the identity and go “around” the group in another direction to get to the same point. Some of the latter may even be shorter. The simplest example of this occurs in $\mathrm{SU}(2)$, whose manifold is a 3-sphere. Geodesics on the 3-sphere are great circles. If we continue on a great circle past the point diametrically opposite to the initial one, there is a shorter geodesic that goes the other way around the sphere. For general $\mathrm{SU}(N)$ there may be many more distinct paths between two group elements that exploit the complicated topology, and finding them all is an unsolved problem. The problem is even believed to be NP-complete from an algorithmic perspective. While finding all global shortcuts on $\mathrm{SU}(N)$ equipped with a given complexity metric is intractable, the authors of [112, 114] pointed out that there is a class of shortcuts that we can analyze, associated to conjugate points along a geodesic. Conjugate points imply that the geodesic we started with is a saddlepoint, not a minimum of the length, so that we can find a shorter path by descending from the saddle. Below we will explain how to find conjugate points, and present physical criteria that delay their occurrence until exponential times under Hamiltonian evolution.
@@
Our goal is to find geodesics on a Lie group manifold in a metric that penalizes “hard” generators that represent infinitesimal transformations that are difficult to implement. First, consider the group generators that span the tangent space to the group manifold at the identity. Consider the Cartan-Killing form $K_{IJ}\propto f_{IM}^{L}f_{JL}^{M}$ where $f_{IJ}^{L}$ are the group’s structure constants defined using the commutation relation for the group generators $[T_{I},T_{J}]=if_{IJ}^{L}T_{L}$ with repeated indices summed. Now define
@@
$G_{IJ}=\begin{pmatrix}\delta_{\alpha\beta}&0\\ 0&(1+\mu)\,\delta_{\dot{\alpha}\dot{\beta}}\end{pmatrix}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ ;\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mu\sim e^{\epsilon S}\,.$ (27)
@@
With these definitions, the geodesic equation on $\mathrm{SU}(N)$ in Euler-Arnold form is [112] [^13] We relabeled the time $t$ along the circuit as $s$ to avoid confusion with the physical time of the system.
@@
This is a practical formalism for calculations. For example, consider the simple case of $\mathrm{SU}(2)$ with generators $T_{1,2}=\gamma_{1,2}$ and $T_{3}=i\gamma_{1}\gamma_{2}$, where $\gamma_{i}$ are the standard Gamma matrices and $T_{3}$ is deemed “hard”, while $T_{1,2}$ are easy [112]. Geodesics take the explicit form
$V^{1}(s)=v^{1}\cos(v^{3}\mu s)-v^{2}\sin(v^{3}\mu s)\leavevmode\nobreak\ \leavevmode\nobreak\ ;\leavevmode\nobreak\ \leavevmode\nobreak\ V^{2}(s)=v^{2}\cos(v^{3}\mu s)+v^{1}\sin(v^{3}\mu s)\leavevmode\nobreak\ \leavevmode\nobreak\ ;\leavevmode\nobreak\ \leavevmode\nobreak\ V^{3}(s)=\frac{v^{3}}{2}\,,$ (31)
@@
Now consider $\mathrm{SU}(N)$ for general $N$. Following [112, 114] we will consider a representation of the group based on the algebra of Gamma matrices $\gamma_{a}$ with $a=0,\dots,N-1$ and $\{\gamma_{a},\gamma_{b}\}=2\delta_{ab}$. We can construct the $2^{N}$ generators of $\mathrm{SU}(N)$ as $T_{a_{1}\cdots a_{m}}=\gamma_{a_{1}}\cdots\gamma_{a_{m}}$ subject to the condition $a_{p}<a_{q}$ if $p<q$. We will label the generators as $T_{I}$ with a multi-index $I\equiv(a_{1}\cdots a_{m})$. We will later discuss a physical realization of the SYK model, constructed from $N$ Majorana fermions $\psi_{a}$ where $\gamma_{a}\sim\psi_{a}$. In view of this identification, we will say that generators $T_{\alpha}$ with no more than $k$ $\gamma$ matrices in their definition are k-local, and that generators $T_{\dot{\alpha}}$ with more than $k$ $\gamma$ matrices in their definition are k-nonlocal. We will also say that k-local generators are “easy” and that k-nonlocal generators are “hard” and are therefore penalized by $1+\mu$ in the metric in (27). Now suppose that the Hamiltonian is k-local: $H=\sum_{\alpha}J^{\alpha}T_{\alpha}$. Then the authors of [112, 114] showed from the Euler-Arnold equation (28) that there is always a geodesic from the identity to $e^{-iHt}$ along the path taken by the physical time evolution. In the notation of (28), the components of the tangent vectors along this geodesic are $V^{\alpha}=J^{\alpha}t$ leading to the tangent vector $V=Ht$. This means that the Nielsen complexity grows linearly in time,

### v3

@@ 5.2 Geometric features of Nielsen complexity
Returning to the cost function (25), consider taking $p=1$. When $p=1$, $F_{p{=}1,\vec{q}}[Y^{I}]$ has a natural physical interpretation: it counts (with an appropriate measure) the number of gates used to build the continuous trajectory in the group manifold SU($N$). However, $F_{p{=}1,\vec{q}}[Y^{I}]$ has the disadvantage of not being smooth. We therefore cannot apply the calculus of variations to the geodesics. When $p=2$, the norm (25) induces a Riemannian metric on the group manifold, simplifying the study of the geodesics. In this case, the map (22) is a Riemannian submersion Auzzi:2020idm (defined in petersen2006riemannian). Heuristically, a Riemannian submersion is a smooth mapping from a higher-dimensional manifold to a lower-dimensional one. This mapping preserves distances in the directions perpendicular to the fibers. [^12] The fibers are points in the higher-dimensional space that are mapped to the same point in the lower-dimensional manifold. In our setting, the fibers correspond to the maximal subgroup $\mathrm{SU}(N-1)\times\mathrm{U}(1)$. The submersion maps $\mathrm{SU}(N)$ to $\mathbb{CP}^{N-1}$. These facts provide a systematic means of determining the metric induced by the cost-function minimization (23) on the state space. The Nielsen complexity literature has mainly focused on $p=1,2$ Jefferson:2017sdb, Chapman:2017rqy, Khan:2018rzm, Hackl:2018ptj, Chapman:2018hou, Bhattacharyya:2018bbv, Guo:2018kzl, Brown:2019whu, Bernamonti:2019zyy, Caceres:2019pgf, Brown:2017jil, Caputa:2018kdj, Balasubramanian:2018hsu, Balasubramanian:2019wgd, Auzzi:2020idm, Chagnet:2021uvi, Basteiro:2021ene, Brown:2021uov, Balasubramanian:2021mxo, Brown:2022phc, Erdmenger:2022lov, Baiguera:2023bhm.
One may wonder whether Nielsen’s complexity is related to discrete gate complexity. The answer is yes: Nielsen’s complexity provides both upper and lower bounds on the gate complexity, provided the cost function satisfies certain conditions Nielsen1, Nielsen2. The authors of Nielsen1, Nielsen2 further show that, by sufficiently penalizing the generators acting on several qubits, one can force the geodesic trajectory that reaches a unitary in the complexity geometry with arbitrary precision to only use generators acting on one or two qubits. The penalties required for this construction typically make the manifold highly curved. However, as we will discuss, Nielsen’s complexity is universal at long distances within a certain equivalence class of metrics, and this equivalence class can be shown to also contain moderately-curved metrics. A priori, this fact makes the computation of Nielsen’s complexity easier by using tools from differential geometry. Finally, let us mention that when the motion in the unitary manifold is restricted to a small subgroup, exact or semi-analytic results can be obtained, which we discuss in later subsections.
@@
There is substantial freedom in choosing a cost function. However, there is a large universality class of cost functions leading to metrics on the group manifold that may vary significantly at short distances, but differ at large distances only polynomially in the path length and in the physical system’s size Brown:2021uov, Brown:2022phc. To simplify the rest of this subsection, we will take the penalty factors to only depend on the number of qubits $k$ on which a gate acts, and denote them as $q_{k}$. Sometimes $k$ is referred to as the weight of the gate – $q_{k}$ should not be confused with the $q_{I}$ introduced in Eq. (25). The notation $q_{k}$ means that generators $T_{I}$ that have the same weight $k$ share the same penalty factor.
Consider increasing any penalty factor while keeping the others constant. Above a critical value $\bar{q}_{k}$, the complexity ceases to depend on that penalty factor because we can implement this generator with weight $k$ as a composition of other, cheaper gates. The reason is as follows. Consider any fixed $m\geq 2$. Using the $m$-local gates, one can reconstruct all the $m^{\prime}$-local gates for every $m^{\prime}$ nielsen2010quantum. Consequently, one can identify a certain critical schedule $\{\bar{q}_{k}\}$ in terms of which we can define a universality class of cost functions that yield metrics with the same long distance properties Brown:2021uov. Specifically, given the critical schedule, we consider penalty factors $\{q_{k}\}$ satisfying $q_{2}=\bar{q}_{2}$ and, for all $m>2$, $q_{m}\geq\bar{q}_{m}$. Metrics arising from this universality class of cost functions may differ greatly at short separations but lead to approximately equal distances at long separations. That is, short and long distance scales decouple. The critical schedule is conjectured to be the only universality class member for which the conjugate points are pushed to infinity, and geodesics are straight lines for an exponential time in the number of qubits Brown:2021uov. [^13] A conjugate point occurs along a geodesic if there is a local shortcut from the start of the trajectory to some point along it (the notion of local shortcut will be more precisely defined below). Such conjugate points can be diagnosed by locally perturbing a geodesic. If the perturbation does not change the length, the geodesic can be deformed to find a shorter geodesic between the same endpoints. There can also be global obstructions to minimality of a geodesic, such as geodesics encircling a sphere in opposite directions along a diameter. We will discuss conjugate points in more detail in Sec. 5.4. Brown argued that the same universality class consists of metrics whose associated Nielsen complexity, measured by the geodesic length from the origin, is polynomially equivalent to the gate complexity. At the same time, it is possible to pick a metric inside the universality class such that the unitary manifold has modest curvature Brown:2022phc, thus making computations easier.
@@ 5.3 Constraints on the cost function from expected time dynamics
In Sec. 4.1, we showed that quantum complexity for discrete time circuits exhibits two behaviors: linear growth which can last for an exponentially long time, and the switchback effect. These derivations relied on simple counting arguments. Afterwards, we quantitatively demonstrated this linear growth for random unitary circuits, by using algebraic geometry and unitary designs (Sections 4.34.6). We would like to understand which complexity geometries reproduce the above key features of complexity when applied to the unitary time evolution generated by a chaotic Hamiltonian. To do so, we must select an appropriate metric on the $\mathrm{SU}(N)$ manifold. The requirements turn out to be Nielsen3, Brown:2016wib, Brown:2017jil:
@@
    The group manifold must have a negative average sectional curvature Nielsen3, Brown:2016wib, such that nearby geodesics diverge. This is necessary, but not sufficient, for establishing the above key features for chaotic and ergodic Hamiltonians.
@@
Let us comment on the implications of these requirements. First, consider Eq. (23) which yields a metric on the space of states rather than on unitaries. Do the above properties have a reflection in the induced geometry on the space of states? In fact the induced metric is not homogeneous, complicating the study of the geodesics and curvature in the space of states Brown:2019whu. Nonetheless, as mentioned earlier, for the case of Riemannian cost functions, the map (22) is a Riemannian submersion Auzzi:2020idm (defined in petersen2006riemannian). Riemannian submersions relate sectional curvatures in the space of unitaries and sectional curvatures in the space of states, via the O’Neill formula o1966fundamental. The latter implies that the sectional curvature along a plane in the space of states is lower-bounded by the sectional curvature along an appropriate plane in the unitary manifold.
@@
Second, the group manifold equipped with the Cartan-Killing metric—the standard metric on SU(N)(N)—does not satisfy the above requirements because it has only constant positive sectional curvatures. How can we construct a geometry that has negative sectional curvatures? We can assign the penalty factors such that some of the $\mathfrak{su}(N)$ algebra’s nonvanishing commutators have the structure [easy,easy]=hard[easy,\,easy]=\,hard. The easy and hard refer to generators associated with small and large penalty factors in the cost function. When the penalty factors are associated to generators such that the commutators take the above-mentioned form, then the Pythagorean theorem for curved space guarantees some negative sectional curvatures Brown:2019whu. The latter statement was proven in the contexts of the unitary manifolds of qubits and qutrits Brown:2019whu, Auzzi:2020idm. [^14] A qudit is a $d$-level quantum system, represented by a Hilbert space $\mathbb{C}^{d}$. A qubit has $d=2$; and a qutrit, $d=3$. Note that manifolds that are equipped with right-invariant metrics, like the ones we are interested in, must also have some positive sectional curvatures; otherwise, they are flat MILNOR1976293.
@@
References Balasubramanian:2019wgd, Balasubramanian:2021mxo studied time-dependent trajectories generated by the SYK Hamiltonian on the unitary manifold. The authors showed that the geodesic distance from the origin to the time evolution operator is lower-bounded by a geodesic whose length grows linearly with time. This linear growth is truncated at conjugate points along the geodesic. There is a shorter path connecting the origin to a conjugate point than the one produced by extending the earlier time geodesic. It is possible that the linear growth is truncated even earlier by global obstructions – geodesics that wind in some other direction around the unitary group to the same point – but these are much more difficult to study, as we will discuss in Sec. 5.4. Conjugate points in the critical complexity geometry described at the end of Sec. 5.2 are pushed to infinity Brown:2021uov. Therefore, the linear growth of Nielsen’s complexity in that geometry persists for a time exponential in the number of qubits. In the context of spin chains, Brown showed that many complexity geometries have diameters that are exponentially large in the number of qubits Brown:2022phc. This exponential relation is necessary for complexity to saturate at an exponentially large value (Fig. 3).
@@
We have argued that Nielsen complexity exhibits one behavior characteristic of quantum complexity: the initial linear growth. Nielsen complexity also reproduces the other behavior—the switchback effect—in toy models Brown:2017jil. Qualitative features similar to the switchback effect are also displayed by single-qubit and two-qubit systems Caginalp:2020tzw. However, a comprehensive picture of the emergence of the switchback effect from the complexity geometry is still lacking and is an open question for future research.
@@ 5.4 Physical criteria for complexity growth: general results and applications
How is Nielsen complexity related to standard concepts in physics, like energy spectra, correlation functions, and Schrödinger dynamics? Several connections were uncovered by the authors of Balasubramanian:2019wgd, Balasubramanian:2021mxo, who investigated physical criteria that control the length of time over which Nielsen complexity will grow during Hamiltonian time evolution. We will discuss their results below, and use this discussion as a way of illustrating how to perform practical calculations of Nielsen complexity.
@@
For concreteness, let us consider a time-independent Hamiltonian acting on an NN–dimensional Hilbert space. Time evolution is then governed by the unitary operator $U(t)=e^{-iHt}$. We want to think of this operator as tracing a trajectory on the $\mathrm{SU}(N)$ group manifold from the identity $\mathbbm{1}$ at $t=0$ to $U(t)$. Following Nielsen, we will regard the complexity of $U(t)$ as measured by the length of the shortest geodesic from $\mathbbm{1}$ to $U(t)$ in a certain complexity metric that we will define below.
@@
In fact, finding the shortest geodesic on $\mathrm{SU}(N)$ for general $N$ is a difficult problem in any metric because the group manifold has a complicated fiber bundle topology. So, if we find one geodesic between two group elements by solving the appropriate differential equation, there may well be others that start at the identity and go “around” the group in another direction to get to the same point. Some of the latter may even be shorter. The simplest example of this occurs in $\mathrm{SU}(2)$, whose manifold is a 3-sphere. Geodesics on the 3-sphere are great circles. If we continue on a great circle past the point diametrically opposite to the initial one, there is a shorter geodesic that goes the other way around the sphere. For general $\mathrm{SU}(N)$ there may be many more distinct paths between two group elements that exploit the complicated topology, and finding them all is an unsolved problem. The problem is even believed to be NP-complete from an algorithmic perspective. While finding all global shortcuts on $\mathrm{SU}(N)$ equipped with a given complexity metric is intractable, the authors of Balasubramanian:2019wgd, Balasubramanian:2021mxo pointed out that there is a class of shortcuts that we can analyze, associated to conjugate points along a geodesic. Conjugate points imply that the geodesic we started with is a saddlepoint, not a minimum of the length, so that we can find a shorter path by descending from the saddle. Below we will explain how to find conjugate points, and present physical criteria that delay their occurrence until exponential times under Hamiltonian evolution.
@@
Our goal is to find geodesics on a Lie group manifold in a metric that penalizes “hard” generators that represent infinitesimal transformations that are difficult to implement. First, consider the group generators that span the tangent space to the group manifold at the identity. Consider the Cartan-Killing form $K_{IJ}\propto f_{IM}^{L}f_{JL}^{M}$ where $f_{IJ}^{L}$ are the group’s structure constants defined using the commutation relation for the group generators $T_{I},T_{J}]=if_{IJ}^{L}T_{L}$ with repeated indices summed. Now define
@@
$G_{IJ}=\begin{pmatrix}\delta_{\alpha\beta}&0\\ 0&(1+\mu)\,\delta_{\dot{\alpha}\dot{\beta}}\end{pmatrix}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ ;\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \mu\sim e^{\epsilon S}\,.$ (27)
@@
With these definitions, the geodesic equation on $\mathrm{SU}(N)$ in Euler-Arnold form is Balasubramanian:2019wgd [^15] We relabeled the time $t$ along the circuit as $s$ to avoid confusion with the physical time of the system.
@@
This is a practical formalism for calculations. For example, consider the simple case of $\mathrm{SU}(2)$ with generators $T_{1,2}=\gamma_{1,2}$ and $T_{3}=i\gamma_{1}\gamma_{2}$, where $\gamma_{i}$ are the standard Gamma matrices and $T_{3}$ is deemed “hard”, while $T_{1,2}$ are easy Balasubramanian:2019wgd. Geodesics take the explicit form
$V^{1}(s)=v^{1}\cos(v^{3}\mu s)-v^{2}\sin(v^{3}\mu s)\penalty 10000\ \penalty 10000\ ;\penalty 10000\ \penalty 10000\ V^{2}(s)=v^{2}\cos(v^{3}\mu s)+v^{1}\sin(v^{3}\mu s)\penalty 10000\ \penalty 10000\ ;\penalty 10000\ \penalty 10000\ V^{3}(s)=\frac{v^{3}}{2}\,$, (31)
@@
Now consider $\mathrm{SU}(N)$ for general $N$. Following Balasubramanian:2019wgd, Balasubramanian:2021mxo we will consider a representation of the group based on the algebra of Gamma matrices $\gamma_{a}$ with $a=0,\dots,N-1$ and $\{\gamma_{a},\gamma_{b}\}=2\delta_{ab}$. We can construct the $2^{N}$ generators of $\mathrm{SU}(N)$ as $T_{a_{1}\cdots a_{m}}=\gamma_{a_{1}}\cdots\gamma_{a_{m}}$ subject to the condition $a_{p}<a_{q}$ if $p<q$. We will label the generators as $T_{I}$ with a multi-index $I\equiv(a_{1}\cdots a_{m})$. We will later discuss a physical realization of the SYK model, constructed from $N$ Majorana fermions $\psi_{a}$ where $\gamma_{a}\sim\psi_{a}$. In view of this identification, we will say that generators $T_{\alpha}$ with no more than $k$ $\gamma$ matrices in their definition are k-local, and that generators $T_{\dot{\alpha}}$ with more than $k$ $\gamma$ matrices in their definition are k-nonlocal. We will also say that k-local generators are “easy” and that k-nonlocal generators are “hard” and are therefore penalized by $1+\mu$ in the metric in (27). Now suppose that the Hamiltonian is k-local: $H=\sum_{\alpha}J^{\alpha}T_{\alpha}$. Then the authors of Balasubramanian:2019wgd, Balasubramanian:2021mxo showed from the Euler-Arnold equation (28) that there is always a geodesic from the identity to $e^{-iHt}$ along the path taken by the physical time evolution. In the notation of (28), the components of the tangent vectors along this geodesic are $V^{\alpha}=J^{\alpha}t$ leading to the tangent vector $V=Ht$. This means that the Nielsen complexity grows linearly in time,